A baseball pitcher throws a 141 g baseball at a velocity of 33 m/s towards home plate. The batter hits the ball, and the velocity of the ball coming off the bat is 54 m/s in the opposite direction of the velocity of the pitch. The bat is in contact with the ball for 0.003 seconds. What is the magnitude in Newtons of the average force applied to the ball?
Answer #1

given m=141g=141*10-3 kg

u=33 m/s

v = -54 m/s

  \small \Delta t=0.003 s

impulse=F\small \Delta t = \small m\Delta v

I=\small m\Delta v = (141*10-3)*(33-(-54))=12.267 kg m2

F\small \Delta t=12.267

F=12.267/0.003= 4089N